# DyCoke[Uncaptioned image]: Dynamic Compression of Tokens for Fast Video Large Language Models

### v1

@@ 4.2 Main Results
Video QA. (1) As shown in Tab. 1, our method significantly outperforms the counterpart methods FastV and PruMerge at similar or lower computational cost, on different benchmarks on 0.5B, 7B and 72B VLLMs.
@@ 4.3 Efficiency Analysis
Latency and Memory Comparison. We first compare the speed and memory consumption of model inference with different numbers of sampling frames (16 and 32 frames). To ensure result robustness, we test on the MVBench dataset [19] to minimize the influence of output length. Results in Tab. 3 show that the model with compressed tokens by our method runs significantly faster than its full-token counerpart, with a speedup of 1.4$\times$ on 7B models. The advantage is even more pronounced for longer visual sequences, with a speedup of 1.54$\times$. The speedup comes along with lower memory consumption than other baseline methods.
@@ 4.4 Ablation Study
Token Selection Strategy. In our proposed TTM (token temporal merging) module, token similarity between adjacent needles and preceding frames is used to evaluate temporal redundancy for pruning and merging. In Tab. 5 (marked in blue background), the effect of random token selection for pruning in TTM on model performance is analyzed. The effectiveness of the visibility model in understanding video content drops substantially, highlighting the superiority of our token filtering strategy. Furthermore, we investigate the effects of overpruning in the TTM module (highlighted in gray) and observe a sharp performance drop, further demonstrating the TTM module’s effectiveness in mitigating the temporal redundancy of visual tokens.
@@ 5 Conclusion
This paper presents DyCoke, a new training-free method to dynamically reduce the visual tokens for faster video large language models (VLLMs). We develop a two-stage token compression strategy that leverages the temporal and spatial redundancy in video information: In the first stage, highly similar temporal tokens among frames are merged; the second stage further reduces the visual tokens used for attention computation during decoding stage. To the best of our knowledge, this is the first dynamic token pruning method specifically tailored for VLLMs. Extensive benchmark and analysis results on a wide range of video QA and reasoning tasks with three VLLMs (0.5B, 7B, 70B parameters) show our method consistently surpasses the prior SoTA counterparts. Using our method on VLLMs, we can achieve up to 1.4$\times$ memory reduction, 1.5$\times$ speedup, with performance still improved.
@@ A.1 Brief Overview
To complement the illustration, we provide a brief description of the 20 tasks included in the MVBench dataset. The MVBench dataset focuses on evaluating the model’s temporal reasoning ability, spanning basic perceptual to advanced cognitive tasks across nine broad categories, including complex tasks such as action recognition, object localization, and scene transformation. Each task requires the model to handle dynamic changes in video sequences, compensating for the limitations in temporal understanding found in existing still-image tasks. For example, in the “action” task, the model must recognize action sequences, predict future actions, and make fine distinctions between similar actions to achieve a nuanced understanding of human behavior in videos. Additionally, MVBench includes tasks involving object interaction and state changes, such as determining whether an object is present in a video or identifying object position changes over different time periods. The dataset also includes high-level cognitive tasks such as “counterfactual reasoning” and “episodic reasoning,” requiring the model to speculate on causality in complex situations and navigate based on an egocentric perspective.
The 20 tasks in the Tab. 2 are: AS (action sequence), AP (action prediction), AA (action antonymy), FA (fine-grained action), UA (unexpected action), OE (object existence), OI (object interaction), OS (object shuffle), MD (movement direction), AL (action localization), ST (scene transition), AC (action counting), MC (movement counting), MA (movement attributes), SC (state change), FP (fine-grained pose), CO (character order), EN (egocentric navigation), ER (episodic reasoning), and CI (counterfactual inference).
@@ Appendix B Model Hyperparameters
In Sec. 4.1, we evaluated token compression using computational cost FLOPs, calculating that multi-head attention (MHA) and feedforward network (FFN) modules are the two primary computational costs. Here, $n$ represents the number of tokens, $d$ is the hidden state size, and $m$ is the intermediate size of the FFN. For the three sizes of VLLMs used in this work, we provide supplementary explanations for $n$, $m$, $d$, and the total number of transformer layers $T$, as shown in Tab. 8.
@@
Appendix C More Visualizations

### v3

@@ 4.2 Main Results
Video QA. (1) As shown in Tab. 1, our method significantly outperforms the counterpart methods FastV and PruMerge at similar or lower computational cost, on different benchmarks on 0.5B, 7B, and 72B VLLMs.
@@ 4.3 Efficiency Analysis
Latency and Memory Comparison. We first compare the speed and memory consumption of model inference with different numbers of sampling frames (16 and 32 frames). To ensure result robustness, we test on the MVBench dataset [23] to minimize the influence of output length. Results in Tab. 3 show that the model with compressed tokens by our method runs significantly faster than its full-token counterpart, with a speedup of 1.4$\times$ on 7B models. The advantage is even more pronounced for longer visual sequences, with a speedup of 1.54$\times$. The speedup comes along with lower memory consumption than other baseline methods.
@@ 4.4 Ablation Study
Token Selection Strategy. In our proposed TTM (token temporal merging) module, the token similarity between adjacent needles and preceding frames is used to evaluate temporal redundancy for pruning and merging. In Tab. 5 (marked in blue background), the effect of random token selection for pruning in TTM on model performance is analyzed. The effectiveness of the visibility model in understanding video content drops substantially, highlighting the superiority of our token filtering strategy. Furthermore, we investigate the effects of over-pruning in the TTM module (highlighted in gray) and observe a sharp performance drop, further demonstrating the TTM module’s effectiveness in mitigating the temporal redundancy of visual tokens.
@@ 5 Conclusion
This paper presents DyCoke, a new training-free method to dynamically reduce the visual tokens for faster video large language models (VLLMs). We develop a two-stage token compression strategy that leverages the temporal and spatial redundancy in video information: In the first stage, highly similar temporal tokens among frames are merged; the second stage further reduces the visual tokens used for attention computation during the decoding stage. To the best of our knowledge, this is the first dynamic token pruning method specifically tailored for VLLMs. Extensive benchmark and analysis results on a wide range of video QA and reasoning tasks with three VLLMs (0.5B, 7B, 70B parameters) show our method consistently surpasses the prior SoTA counterparts. Using our method on VLLMs, we can achieve up to 1.4$\times$ memory reduction, and 1.5$\times$ speedup, with performance still improved.
@@ A.1 Brief Overview
To complement the illustration, we provide a brief description of the 20 tasks included in the MVBench dataset. The MVBench dataset focuses on evaluating the model’s temporal reasoning ability, spanning basic perceptual to advanced cognitive tasks across nine broad categories, including complex tasks such as action recognition, object localization, and scene transformation. Each task requires the model to handle dynamic changes in video sequences, compensating for the limitations in temporal understanding found in existing still-image tasks. For example, in the “action” task, the model must recognize action sequences, predict future actions, and distinguish between similar actions to achieve a nuanced understanding of human behavior in videos. Additionally, MVBench includes tasks involving object interaction and state changes, such as determining whether an object is present in a video or identifying object position changes over different periods. The dataset also includes high-level cognitive tasks such as “counterfactual reasoning” and “episodic reasoning,” requiring the model to speculate on causality in complex situations and navigate based on an egocentric perspective. The 20 tasks in the Tab. 2 are: AS (action sequence), AP (action prediction), AA (action antonymy), FA (fine-grained action), UA (unexpected action), OE (object existence), OI (object interaction), OS (object shuffle), MD (movement direction), AL (action localization), ST (scene transition), AC (action counting), MC (movement counting), MA (movement attributes), SC (state change), FP (fine-grained pose), CO (character order), EN (egocentric navigation), ER (episodic reasoning), and CI (counterfactual inference).
@@ Appendix B Model Hyperparameters
In Sec. 4.1, we evaluated token compression using computational cost FLOPs, calculating that multi-head attention (MHA) and feedforward network (FFN) modules are the two primary computational costs. Here, $n$ represents the number of tokens, $d$ is the hidden state size and $m$ is the intermediate size of the FFN. For the three sizes of VLLMs used in this work, we provide supplementary explanations for $n$, $m$, $d$, and the total number of transformer layers $T$, as shown in Tab. 8.
@@
Appendix C Computing Cost Evaluation.
We examine the total FLOPs of the prefilling stage and the decoding stage. Consider a transformer layer employing multi-head attention (MHA) and feed-forward network (FFN) modules. Let $n$, $d$, and $m$ denote the number of tokens, the hidden state size, and the intermediate size of the FFN, respectively. In the prefilling phase, the total FLOPs can be estimated as $4nd^{2}+2n^{2}d+2ndm$. For the decoding phase, considering the significant contribution of the KV cache, the computational consumption for $R$ total iterations (i.e., predicting $R$ tokens) is $R\left(4d^{2}+2dm\right)+2\sum_{i=1}^{R}d\times(n+i)$. We unify $R=100$ for calculation in the experiments. Thus, for an LLM with $T$ total transformer layers, the total FLOPs can be expressed as follows,
$\mathrm{FLOPs}$ $=T(4nd^{2}+2n^{2}d+2ndm)$ (8)
$+TR\left((4d^{2}+2dm)+2\left(dn+\frac{d(R+1)}{2}\right)\right).$
FLOPs are employed as a metric to quantify token computation, ensuring a fair comparison with other methods; however, they do not directly indicate the final inference speed.
Appendix D Ablation Study about $k$ and Input Frames
Figure 5: Performance vs. $K$ values in different input frames.
As shown in Figure 5, we investigate the relationship among the numerical value of $k$, the number of input frames (32, 16, and 8), and the overall model performance. Our results indicate that with a low number of input frames, token compression consistently leads to a decline in model performance. However, as the number of input frames increases—that is, as the multiplicity of visual tokens grows—the adverse impact of token compression on model performance gradually diminishes, eventually outperforming the baseline model. This phenomenon arises because more input frames introduce increased information redundancy and noise, which can be mitigated through moderate token compression, thereby maintaining performance with a slight enhancement.
Appendix E Discussion and Future Work
E.1 Compatible with Flash Attention
Flash Attention requires additional computation during the inference stage to compute the attention score matrix. However, combining Dycoke with Flash Attention does not impose significant additional computational overhead, as the attention score is computed only at a specific layer during each decoding iteration. Moreover, the computational complexity is substantially lower than that of the prefilling phase.
E.2 Future Work
DyCoke marks the first significant advancement in dynamic token pruning to improve inference efficiency in video large language models (VLLMs), yet some challenges remain for further exploration. Firstly, although DyCoke’s compression strategy effectively reduces token redundancy, specific video contexts (e.g., rapid scene changes or critical time shifts) may still incur minor information loss. While the dynamic token selection mechanism mitigates this risk, future work will focus on developing more fine-grained token compression methods for highly dynamic video content. Secondly, although token compression reduces memory consumption and enhances reasoning speed, fully deploying LLMs on mobile devices remains challenging due to their scale. Thus, we aim to integrate advanced compression techniques, such as quantization and distillation, to develop more efficient VLLMs.
Appendix F More Visualizations